Average Rate of Return—New Product Galactic Inc. is considering an investment in new equipment that will be used to manufacture a smartphone. The phone is expected to generate additional annual sales of 4,100 units at $311 per unit. The equipment has a cost of $343,200, residual value of $25,800, and an eight-year life. The equipment can only be used to manufacture the phone. The cost to manufacture the phone follows: Cost per unit: Direct labor $55.00 Direct materials 212.00 Factory overhead (including depreciation) 35.90 Total cost per unit $302.90 Determine the average rate of return on the equipment. If required, round to the nearest whole percent. %
Answer #1

Average rate of return = Average net income / Average investment

Average net income = (311 - 302.90) *4100 = $33,210

Average investment = (343200+25800) /2 = $184,500

Average rate of return = 33210 / 184500 = $18%